OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (anonymous):

Let Tim's age now = x years Then coach Johnson's age now = (x + 10) years Three years ago, Tim's age was = (x - 3) years Coach Johnson's age was = (x + 10 - 3) years = (x + 7) years ATQ, x+ 7 = 2(x - 3) x + 7 = 2x - 6 7 + 6 = 2x - x 13 = x So Tim is now 13 years old.